OpenStudy (anonymous):

what is the slope of x+9y=18?

OpenStudy (anonymous):

i am stuck on this one please help

OpenStudy (anonymous):

set in terms of y.... x+9y=18 is the same as 9y=18-x " " \[y=18/9-1/9 x\]

OpenStudy (anonymous):

so if you remember the form y=mx+b our b= 18/9 = 2 and our m= -1/9 so our slope is -1/9